# The Geometry of Codes for Random Access in DNA Storage

### v1

@@ 4.2 Fundamental points with multiplicity
In this subsection, we improve the construction of the point set considered in Subsection 4.1 by using a balanced quasi-arc of weight $x$ and adding an additional $y-1$ copies of each fundamental point to the set where $x\geq 0$ and $y\geq 1$ are integers, resulting in a point set where each fundamental point has multiplicity $y$. We denote this point set by $\mathscr{G}_{x,y}=\mathscr{F}_{y}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$, where $\mathscr{F}_{y}$ contains $y$ copies of $E_{1}$, $E_{2}$, and $E_{3}$, and $\mathscr{G}_{1}$, $\mathscr{G}_{2}$, and $\mathscr{G}_{3}$ satisfy the properties needed for $\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$ to form a balanced quasi-arc. Adding multiplicities to the fundamental points creates a greater imbalance in this multiset, making it even more likely to recover fundamental points. We proceed here similarly to Subsection 4.1 and start with giving closed formulas for $\alpha_{F}(\mathscr{G}_{x,y},s)$ for all $11\leq s\leq n-$ 1.
@@
As in the proof of Proposition 4.2, $\alpha_{E_{1}}(\mathscr{G}_{x,y},s)$ for any $11\leq s\leq n-$ 1 does not depend on $E_{1}$, and so w.l.o.g. suppose that we want to recover the fundamental point $E_{1}$.
@@
As before, for $s=2$, we can count the number of 2-sets in $\mathscr{G}$ that recover $E_{1}$ by distinguishing between three cases: If the two points span $E_{1}E_{2}$ or $E_{3}E_{1}$, they will for sure recover $E_{1}$, and there are $xy+\binom{x}{2}$ 2-sets for either of the lines without considering that $E_{1}$ could be one of the points. If $E_{1}$ is one of the two points, then any of the remaining $3x+2y$ points will give a recovering 2-set. Finally, if $y\geq 2$, the only 2-set that recovers $E_{1}$ that we did not count yet, is the 2-set made of only $E_{1}$. There are $y(y-1)/2$ such sets.
For $3\leq s\leq x+2y$, the only $s$-sets that do not recover $E_{1}$ are those that only consist only of points in $E_{2}E_{3}$, and those that consist of $(s-1)$-sets of either $E_{2}$ or $E_{3}$, together with the fundamental line that does not go through the $E_{2}$ or $E_{3}$ respectively. In total this gives $\binom{3x+3y}{s}-\binom{x+2y}{s}-2\binom{y}{s-1}x$ such $s$-sets.
@@
For $x=y$ the expression in Proposition 4.7 reads and simplifies as follows.
@@
Recall that $\binom{x}{s-1}=\frac{s}{x+1}\binom{x+1}{s}$; see e.g. [28, Table 174, Equation 3]. Using this we get
@@
From [28, page 173], we then see that $\frac{\binom{x+1}{s}}{\binom{6x-1}{s}}=\frac{\binom{6x-1-s}{x+1-s}}{\binom{6x-1}{x+1}}$ and hence
@@
From [28, Table 174, Equation 8] we have that $(x+1)\sum_{T=0}^{x-2}\binom{5x-2+T}{T}=(x+1)\binom{6x-3}{x-2}=(x+1)(x-1)\frac{\binom{6x-3}{x-1}}{5x-1}$, and hence
@@
From experiments with a computer algebra program, it appears that the optimal ratio of $y/x$ for achieving the smallest $\mathbb{E}[\tau_{F}(\mathscr{G}_{x,y})]$ for large $x$ and $y$ is close to $y/x=0.85$. However, it is not obvious to us how to prove this. Nevertheless, we include a plot in Figure 1 that illustrates this behavior for various values of $x$ and different ratios $y/x$.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.41.61.61.61.61.81.81.81.822220.880.880.880.880.90.90.90.90.920.920.920.920.940.940.940.94Ratio $y/x$$x=5$$x=10$$x=100$$x=1000$$x=5000$
@@ 5 On a rate $1/2$ construction for any dimension
Inspired by Theorem 5.1, in order to construct “good” codes for the random access problem, one could try to construct sets of points $\mathscr{G}$ for which the fundamental points are balanced, and the non-fundamental points are balanced, and for which ${\alpha}_{F}(\mathscr{G},s)\geq{\alpha}_{N}(\mathscr{G},s)$ for all $11\leq s\leq n-$ 1 and ${\alpha}_{F}(\mathscr{G},s)>{\alpha}_{N}(\mathscr{G},s)$ where the inequality is strict for some values of $s$. Then, since on average, recovering any of the $n$ points is $k$, and it is “easier” to recover fundamental points than non-fundamental points, clearly the random access expectation for fundamental points has to be smaller than $k$. This idea is formalized and proved in the following lemma, which is one of the main tools for proving the conjecture.
@@
where in the first equality we used the fact that $\alpha_{F}(\mathscr{G},s)=\alpha_{N}(\mathscr{G},s)$ for $n-d+1\leq s\leq n-1$, which holds because a code of minimum distance $d$ can correct $n-d+1$ erasures, and so any $n-d+1$ or more points in $\mathscr{G}$ will recover the remaining points. Combining this with the statement of Theorem 5.1 we obtain
@@
Let $E_{i}$ denote the $i$-th fundamental point in $\mathrm{PG}(k-1,q)$ and let $E_{i,j}$ denote the point corresponding to the sum $E_{i}+E_{j}$. We consider the point set $\mathscr{G}=\mathscr{F}\cup\mathscr{N}$ where =$\mathscr{F}=\{E_{1},\dots,E_{k}\}$ is the set of all fundamental points in $\mathrm{PG}(k-1,q)$, and $\mathscr{N}=\{E_{1,2},E_{2,3},\dots,E_{k-1,k},E_{k,1}\}$ is a set of non-fundamental points obtained by summing consecutive fundamental points.
@@
In the sequel, we denote $\beta_{F}(\mathscr{G}_{k},s):=\binom{2k}{s}-\alpha_{F}(\mathscr{G}_{k},s)$ and $\beta_{N}(\mathscr{G}_{k},s):=\binom{2k}{s}-\alpha_{N}(\mathscr{G}_{k},s)$ for all $11\leq s\leq 2k-$ 1 . Note that contrary to $\alpha_{F}(\mathscr{G}_{k},s)$ and $\alpha_{N}(\mathscr{G}_{k},s)$, $\beta_{F}(\mathscr{G}_{k},s)$ and $\beta_{N}(\mathscr{G}_{k},s)$, respectively, count the number of subsets of cardinality $s$ in $\mathscr{G}_{k}$ that do not recover fundamental points, and non-fundamental points, respectively.
@@
We compute a recursive formula for the number of $s$-sets of points in $\mathscr{G}_{k}$ that recover $E_{1}$ by taking the same approach to the one in the proof of [2, Theorem 10]. Consider the directed graph $G_{k}=(V_{k},E_{k})$ whose vertices are made of the $12k-$ 1 points in $\mathscr{G}_{k}\setminus E_{1}$, and whose edges can be partitioned into blue and red edges, which are characterized as follows.
@@
    For every $21\leq i\leq k-$ 2 there are two outgoing blue edges from $E_{i,i+1}$: one to the vertex $E_{i+1}$ and one to the vertex $E_{i+1,i+2}$;
@@
    For every $12\leq i\leq k-$ 1 there are two outgoing red edges from $E_{i,i+1}$: one to the vertex $E_{i}$ and one to the vertex $E_{i-1,i}$;
@@
    Proceeding with the steps above we end up with a total of $\sum_{j=1}^{s}\binom{2k-2j-3}{s-j}$ subgraphs that will not recover $E_{1}$. Finally, note that in the special case where $s=k-1$ we did not count the subgraph that consists of the vertices in $\{E_{2,3},E_{3,4},\dots,E_{k-1,k},E_{k,1}\}$.
    Moreover, when $s\geq 2k-2$, then any subset of size $s$ will recover $E_{1}$, and the other cases of the proposition follow directly from the definitions.
Unlike the case of fundamental points, for the non-fundamental points we were not able to find a recursive formula, but we were able to provide a lower bound on $\beta_{N}(\mathscr{G}_{k},s)$ depending also on $\beta_{N}(\mathscr{G}_{k-1},s-1)$.
@@
$\beta_{N}(\mathscr{G}_{k},s)\geq\begin{cases}0\quad&\textnormal{if $k=1$, $s\geq 1$}\\ 2\quad&\textnormal{if $k=2$, $s=1$}\\ 2k-1\quad&\textnormal{if $k\geq 3$, $s=1$}\\ 0\quad&\textnormal{if $k=2$, $s\geq 2$}\\ 8\quad&\textnormal{if $k=3$, $s=2$}\\ 2\quad&\textnormal{if $k=3$, $s=3$}\\ \binom{2k-3}{s}+\beta_{N}(\mathscr{G}_{k-1},s-1)+\\ \displaystyle\sum_{j=0}^{s-1}\binom{2k-2j-5}{s-j-1}+\displaystyle\sum_{j=0}^{s-2}\binom{2k-2j-7}{s-j-2}+\\ 2\displaystyle\sum_{j=0}^{s-3}\binom{2k-2j-7}{s-j-3}+\displaystyle\sum_{j=0}^{s-4}\binom{2k-2j-7}{s-j-4}\quad&\textnormal{if $k\geq 4$, $2\leq s\leq 2k-3$}\\ 0\quad&\textnormal{if $k\geq 3$, $s\geq 2k-2$}\\ \end{cases}$
@@
We compute a recursive formula by counting the number of $s$-sets of points in $\mathscr{G}_{k}$ that recover $E_{1,2}$. Similarly to the proof of Proposition 5.6, the arguments we use are based on counting certain paths in an appropriate directed graph: consider the directed graph $G_{k}=(V_{k},E_{k})$ whose vertices are made of the $12k-$ 1 points in $\mathscr{G}_{k}\setminus E_{1,2}$, and whose set of edges consists of blue and red edges, defined as follows.
@@
    For every $22\leq i\leq k-$ 2 there are two outgoing blue edges from $E_{i,i+1}$: one to the vertex $E_{i+1}$ and one to the vertex $E_{i+1,i+2}$;
@@
    For every $3\leq i\leq k$ there is an outgoing blue edge to $E_{1}$;
@@
    For every $12\leq i\leq k-$ 1 there are two outgoing red edges from $E_{i,i+1}$: one to the vertex $E_{i}$ and one to the vertex $E_{i-1,i}$;
@@
    For every $3\leq i\leq k$ there is an outgoing red edge to $E_{2}$.
@@
    If $E_{2,3}\notin V$ and $E_{2}\in V$, then $E_{1}\notin V$. Moreover, $V$ cannot contain a red path starting in $E_{k,1}$ and ending in one of the vertices $E_{i}$ for $3\leq i\leq k$. We have counted the number of such subgraphs in part 3. of the proof of Proposition 5.6, and it is equal to $\sum_{j=0}^{s-1}\binom{2k-2j-5}{s-j-1}$. Note that we again have a special case when $s=k-1$, however, since we only give a lower bound for $\beta(\mathscr{G}_{k},s)$, we do not treat this extra case.
@@
    If $E_{2,3},E_{3},E_{3,4}\in V$ then also $E_{1}\notin V$. Using analogous reasoning as before we obtain that the number of such subgraphs is $\sum_{j=0}^{s-3}\binom{2k-2j-7}{s-j-3}$ such paths if $E_{2}\notin V$, and there are $\sum_{j=0}^{s-4}\binom{2k-2j-7}{s-j-4}$.
@@
Since $\alpha_{F}(\mathscr{G}_{k},s)=\binom{2k}{s}-\beta_{F}(\mathscr{G}_{k},s)$ and $\alpha_{N}(\mathscr{G}_{k},s)=\binom{2k}{s}-\beta_{N}(\mathscr{G}_{k},s)$, from the formulas in Propositions 5.6 and 5.7 we obtain that for $k\geq 4$ and $32\leq s\leq 2k-$ 3 we have
@@
Finally, with the asymptotic estimate in Theorem 5.8 we can prove Conjecture 5.5.
@@
Since ${70318847}/{74364290}\approx 0.945599655<0.9456$, Corollary 5.9 proves Conjecture 5.5.

### v2

@@ 4.2 Fundamental points with multiplicity
In this subsection, we improve the construction of the point set considered in Subsection 4.1 by using a balanced quasi-arc of weight $x$ and adding an additional y1y-1 copies of each fundamental point to the set where $x\geq 0$ and $y\geq 1$ are integers, resulting in a point set where each fundamental point has multiplicity $y$. We denote this point set by $\mathscr{G}_{x,y}=\mathscr{F}_{y}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$, where $\mathscr{F}_{y}$ contains $y$ copies of $E_{1}$, $E_{2}$, and $E_{3}$, and $\mathscr{G}_{1}$, $\mathscr{G}_{2}$, and $\mathscr{G}_{3}$ satisfy the properties needed for $\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$ to form a balanced quasi-arc. Adding multiplicities to the fundamental points creates a greater imbalance in this multiset, making it even more likely to recover fundamental points. We proceed here similarly to Subsection 4.1 and start with giving closed formulas for $\alpha_{F}(\mathscr{G}_{x,y},s)$ for all $1\leq s\leq n-1$.
@@
As in the proof of Proposition 4.2, $\alpha_{E_{1}}(\mathscr{G}_{x,y},s)$ for any $1\leq s\leq n-1$ does not depend on $E_{1}$, and so w.l.o.g. suppose that we want to recover the fundamental point $E_{1}$.
@@
As before, for $s=2$, we can count the number of 2-sets in $\mathscr{G}$ that recover $E_{1}$ by distinguishing between three cases: If the two points span $E_{1}E_{2}$ or $E_{3}E_{1}$, they will for sure recover $E_{1}$, and there are $xy+\binom{x}{2}$ 2-sets for either of the lines without considering that $E_{1}$ could be one of the points. If $E_{1}$ is one of the two points, then any of the remaining $3x+2y$ points will give a recovering 2-set. Finally, if $y\geq 2$, the only 2-set that recovers $E_{1}$ that we did not count yet, is the 2-set made of only $E_{1}$. There are y(y1)/2y(y-1)/2 such sets.
For $3\leq s\leq x+2y$, the only $s$-sets that do not recover $E_{1}$ are those that only consist of points in $E_{2}E_{3}$, and those that consist of (s1)(s-1)-sets of either $E_{2}$ or $E_{3}$, together with a point on the fundamental line that does not go through the $E_{2}$ or $E_{3}$, respectively. In total this gives $\binom{3x+3y}{s}-\binom{x+2y}{s}-2\binom{y}{s-1}x$ such $s$-sets.
@@
For $x=y$ the expression in Corollary 4.7 reads and simplifies as follows.
@@
Recall that $\binom{x}{s-1}=\frac{s}{x+1}\binom{x+1}{s}$; see e.g. [30, Table 174, Equation 3]. Using this we get
@@
From [30, page 173], we then see that $\frac{\binom{x+1}{s}}{\binom{6x-1}{s}}=\frac{\binom{6x-1-s}{x+1-s}}{\binom{6x-1}{x+1}}$ and hence
@@
From [30, Table 174, Equation 8] we have that $(x+1)\sum_{T=0}^{x-2}\binom{5x-2+T}{T}=(x+1)\binom{6x-3}{x-2}=(x+1)(x-1)\frac{\binom{6x-3}{x-1}}{5x-1}$, and hence
@@
In this paper we investigate the asymptotic estimate of the formula in Corollary 4.7 for the case where $x=y$. As a follow-up to this work, in [31] the authors provide an asymptotic estimate for $\mathbb{E}[\tau_{F}(\mathscr{G}_{x,y})]$} following the same steps as we took, for the case $x\neq y$, showing that when $y=0.834x$ one obtains $\lim_{x\to\infty}{\mathbb{E}[\tau_{F}(\mathscr{G}_{x,y})]}\leq 0.881\overline{66}\cdot 3$; see [31, Corollary 2]. We include a plot in Figure 1 that illustrates the asymptotics for $\mathbb{E}[\tau_{F}(\mathscr{G}_{x,y})]$ for various values of $x$ and different ratios $y/x$.
00.20.20.40.40.60.60.80.8111.21.21.41.41.61.61.81.8220.880.880.90.90.920.920.940.94Ratio y/xy/xx=5x=5x=10x=10x=100x=100x=1000x=1000x=5000x=5000
@@ 5 On a rate $1/2$ construction for any dimension
Inspired by Theorem 5.1, in order to construct “good” codes for the random access problem, one could try to construct sets of points $\mathscr{G}$ for which the fundamental points are balanced, and the non-fundamental points are balanced, and for which ${\alpha}_{F}(\mathscr{G},s)\geq{\alpha}_{N}(\mathscr{G},s)$ for all $1\leq s\leq n-1$ and ${\alpha}_{F}(\mathscr{G},s)>{\alpha}_{N}(\mathscr{G},s)$ where the inequality is strict for some values of $s$. Then, since on average, recovering any of the $n$ points is $k$, and it is “easier” to recover fundamental points than non-fundamental points, clearly the random access expectation for fundamental points has to be smaller than $k$. This idea is formalized and proved in the following lemma, which is one of the main tools for proving the conjecture.
@@
where in the first equality we used the fact that $\alpha_{F}(\mathscr{G},s)=\alpha_{N}(\mathscr{G},s)$ for $n-d+1\leq s\leq n-1$, which holds because a code of minimum distance $d$ can correct d1d-1 erasures, and so any nd+1n-d+1 or more points in $\mathscr{G}$ will recover the remaining points. Combining this with the statement of Theorem 5.1 we obtain
@@
Let $E_{i}$ denote the $i$-th fundamental point in $\mathrm{PG}(k-1,q)$ and let $E_{i,j}$ denote the point corresponding to the sum $E_{i}+E_{j}$. We consider the point set $\mathscr{G}=\mathscr{F}\cup\mathscr{N}$ where $\mathscr{F}=\{E_{1},\dots,E_{k}\}$ is the set of all fundamental points in $\mathrm{PG}(k-1,q)$, and $\mathscr{N}=\{E_{1,2},E_{2,3},\dots,E_{k-1,k},E_{k,1}\}$ is a set of non-fundamental points obtained by summing consecutive fundamental points.
@@
In the sequel, we denote $\beta_{F}(\mathscr{G}_{k},s):=\binom{2k}{s}-\alpha_{F}(\mathscr{G}_{k},s)$ and $\beta_{N}(\mathscr{G}_{k},s):=\binom{2k}{s}-\alpha_{N}(\mathscr{G}_{k},s)$ for all $1\leq s\leq 2k-1.$ Note that contrary to $\alpha_{F}(\mathscr{G}_{k},s)$ and $\alpha_{N}(\mathscr{G}_{k},s)$, $\beta_{F}(\mathscr{G}_{k},s)$ and $\beta_{N}(\mathscr{G}_{k},s)$, respectively, count the number of subsets of cardinality $s$ in $\mathscr{G}_{k}$ that do not recover fundamental points, and non-fundamental points, respectively.
@@
We compute a recursive formula for the number of $s$-sets of points in $\mathscr{G}_{k}$ that recover $E_{1}$ by taking the same approach to the one in the proof of [2, Theorem 10]. Consider the directed graph $G_{k}$ whose vertices are made of the 2k12k-1 points in $\mathscr{G}_{k}\setminus E_{1}$, and whose edges can be partitioned into blue and red edges, which are characterized as follows.
@@
    For every $1\leq i\leq k-2$ there are two outgoing blue edges from $E_{i,i+1}$: one to the vertex $E_{i+1}$ and one to the vertex $E_{i+1,i+2}$;
@@
    For every $2\leq i\leq k-1$ there are two outgoing red edges from $E_{i,i+1}$: one to the vertex $E_{i}$ and one to the vertex $E_{i-1,i}$;
@@
    Proceeding with the steps above we end up with a total of $\sum_{j=1}^{s}\binom{2k-2j-3}{s-j}$ subgraphs that will not recover $E_{1}$. Finally, note that in the special case where s=k1s=k-1 we did not count the subgraph that consists of the vertices in $\{E_{2,3},E_{3,4},\dots,E_{k-1,k},E_{k,1}\}$.
    Moreover, when $s\geq 2k-2$, then any subset of size $s$ will recover $E_{1}$, and the other cases of the proposition follow directly from the definitions. ∎
Unlike the case of fundamental points, for the non-fundamental points we were not able to find a recursive formula, but we were able to provide a lower bound on $\beta_{N}(\mathscr{G}_{k},s)$ depending also on $\beta_{N}(\mathscr{G}_{k-1},s-1)$. Note it is very likely that one can find an exact recursive formula for $\beta_{N}(\mathscr{G}_{k},s)$, but this would require a more detailed case analysis. For the purpose of this paper, however, our lower bound suffices.
@@
[…]
@@
We compute a recursive formula by counting the number of $s$-sets of points in $\mathscr{G}_{k}$ that recover $E_{1,2}$. Similarly to the proof of Proposition 5.6, the arguments we use are based on counting certain paths in an appropriate directed graph: consider the directed graph $G_{k}$ whose vertices are made of the 2k12k-1 points in $\mathscr{G}_{k}\setminus E_{1,2}$, and whose set of edges consists of blue and red edges, defined as follows.
@@
    For every $2\leq i\leq k-2$ there are two outgoing blue edges from $E_{i,i+1}$: one to the vertex $E_{i+1}$ and one to the vertex $E_{i+1,i+2}$;
@@
    For every $3\leq i\leq k$ there is an outgoing blue edge from $E_{i}$ to $E_{1}$;
@@
    For every $2\leq i\leq k-1$ there are two outgoing red edges from $E_{i,i+1}$: one to the vertex $E_{i}$ and one to the vertex $E_{i-1,i}$;
@@
    For every $3\leq i\leq k$ there is an outgoing red edge from $E_{i}$ to $E_{2}$.
@@
    If $E_{2,3}\notin V$ and $E_{2}\in V$, then $E_{1}\notin V$. Moreover, $V$ cannot contain a red path starting in $E_{k,1}$ and ending in one of the vertices $E_{i}$ for $3\leq i\leq k$. We have counted the number of such subgraphs in part 3. of the proof of Proposition 5.6, and it is equal to $\sum_{j=0}^{s-1}\binom{2k-2j-5}{s-j-1}$. Note that we again have a special case when s=k1s=k-1, however, since we only give a lower bound for $\beta(\mathscr{G}_{k},s)$, we do not treat this extra case.
@@
    If $E_{2,3},E_{3},E_{3,4}\in V$ then also $E_{1}\notin V$. Using analogous reasoning as before we obtain that the number of such subgraphs is $\sum_{j=0}^{s-3}\binom{2k-2j-7}{s-j-3}$ such paths if $E_{2}\notin V$, and there are $\sum_{j=0}^{s-4}\binom{2k-2j-7}{s-j-4}$ if $E_{2}\in V$.
@@
Since $\alpha_{F}(\mathscr{G}_{k},s)=\binom{2k}{s}-\beta_{F}(\mathscr{G}_{k},s)$ and $\alpha_{N}(\mathscr{G}_{k},s)=\binom{2k}{s}-\beta_{N}(\mathscr{G}_{k},s)$, from the formulas in Propositions 5.6 and 5.7 we obtain that for $k\geq 4$ and $2\leq s\leq 2k-3$ we have
@@
Finally, with the asymptotic estimate in Theorem 5.8 we can prove Conjecture $5.$
@@
Since ${70318847}/{74364290}\approx 0.945599655<0.9456$, Corollary 5.9 proves Conjecture $5.$